# “It Explains What I am Currently Going Through Perfectly to a Tee”: Understanding User Perceptions on LLM-Enhanced Narrative Interventions

### v1

@@ Abstract.
Stories about overcoming personal struggles can effectively illustrate the application of psychological theories in real life, yet they may fail to resonate with individuals’ experiences. In this work, we employ large language models (LLMs) to create tailored narratives that acknowledge and address unique challenging thoughts and situations faced by individuals. Our study, involving 346 young adults across two settings, demonstrates that LLM-enhanced stories were perceived to be better than human-written ones in conveying key takeaways, promoting reflection, and reducing belief in negative thoughts. These stories were not only seen as more relatable but also similarly authentic to human-written ones, highlighting the potential of LLMs in helping young adults manage their struggles. The findings of this work provide crucial design considerations for future narrative-based digital mental health interventions, such as the need to maintain relatability without veering into implausibility and refining the wording and tone of AI-enhanced content.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Stories about individuals overcoming personal struggles can serve as a powerful medium for promoting wellbeing (Bhattacharjee et al., 2022; Anderson and Wallace, 2015; Llewellyn-Beardsley et al., 2019; Kornfield et al., 2022b). By engaging with narratives that mirror their own experiences, individuals can draw parallels between their lives and those of the characters, fostering a space for self-reflection (Ferguson and Musheno, 2000). For instance, a narrative about someone overcoming disappointment from job search rejection could provide valuable insights to others facing similar setbacks. This reflection process can help readers recognize patterns in their behavior or thinking, leading to practical revelations that empower them to make positive changes (Bhattacharjee et al., 2023a, 2022). Moreover, engaging with such stories can validate and normalize the mental health challenges that individuals face (Kornfield et al., 2022a).
@@
Given their substantial benefits, stories are increasingly being integrated into digital mental health (DMH) interventions (Meyerhoff et al., 2024; Bhattacharjee et al., 2022; Baumel et al., 2018). However, a significant challenge that many DMH interventions face, including those using storytelling, is their inability to deliver content tailored to the specific challenges that individual users encounter (Matthews and Rhodes-Maquaire, 2024). Many existing DMH interventions cannot be heavily tailored to users’ unique situations, instead relying on pre-defined rules to make coarse content adjustments that may not always resonate (Bhattacharjee et al., 2023b). This misalignment can cause the content to appear generic and disconnected from users’ immediate needs. Such perceptions not only increase the likelihood that the content will be overlooked but also risk dismissing users’ genuine struggles. Consequently, seemingly irrelevant narratives may undermine the potential effectiveness of a DMH intervention (Bhattacharjee et al., 2022).
@@
Large language models (LLMs) are rapidly gaining recognition for their capabilities as personalized intervention tools and supportive scaffolds in a variety of settings (Bhattacharjee et al., 2024b; Kim et al., 2023; Abd-Alrazaq et al., 2023). These models excel in processing open-ended user inputs, enabling them to dynamically adapt and respond to a wide range of queries. Given their flexibility, LLMs could be tools that interpret users’ descriptions of their situations in order to create stories specifically tailored to one’s circumstances. We posit that such capability will significantly enhance the personalization of digital narrative interventions, making content more relevant and impactful. Revisiting the previous example of job search rejection, a user might gain more benefits from a narrative tailored to reflect the competitive nature of their field or environment. Alternatively, the story could address a specific cognitive distortion where they see setbacks in their job search as an ongoing pattern of defeat, failing to recognize their successes or factors out of their control. By connecting the story to specific circumstances and concerns, a narrative might provide deeper insights and more appropriate strategies, effectively addressing the user’s comprehensive experience.
@@
Although LLMs offer these promises for generating personalized stories, implementing them effectively and reliably is non-trivial. They are not innately trained and lack the clinical and experiential knowledge to generate stories on their own (Bhattacharjee et al., 2024b). Considering the sensitivity of mental health struggles, applications of LLMs must avoid generating content that is harmful or insensitive to users’ needs (Demszky et al., 2023). To mitigate these risks, one approach could involve preloading LLMs with narratives and mental health content from previous DMH interventions, limiting them to generate content constrained within those boundaries. This would allow personalization to target changing the context of the story to match an individual’s circumstances while keeping the core takeaways and guidance in line with validated approaches.
In this work, we investigated how users perceive various aspects of LLM-enhanced stories designed to assist individuals in managing negative thoughts. To ensure that LLMs produce appropriate stories, our strategy integrated narratives gathered from real-life experiences that were curated by experts and previously utilized in DMH interventions (Meyerhoff et al., 2024). These stories were accompanied by prompt engineering emphasizing the intended qualities and structure of the stories to be generated. Our aim was to observe how individuals perceive these LLM-enhanced narratives and gather feedback on story qualities, such as the authenticity of the narratives and how effectively key takeaways were communicated. Additionally, we assessed the impact of these stories on changing participants’ beliefs in their negative thoughts and the intensity of their emotions.
@@
Our research focused on young adults aged 18–25, a demographic particularly susceptible to mental health issues due to the inherent transitional nature of this life stage, encompassing significant changes in education, career, and personal relationships (Kornfield et al., 2022a). Statistically, this age group exhibits a much higher prevalence of mental health concerns compared to other adult groups (National Institute of Mental Health (2020), NIMH). These factors make them a relevant group for our research, where we aim to design stories that can help manage common challenges they face.
@@
  • RQ1: How do young adults perceive the key narrative qualities of LLM-enhanced stories, such as authenticity and the ability to communicate takeaways, compared to human-written stories from past DMH interventions?
  • RQ2: How effective are LLM-enhanced stories as interventions for managing negative thoughts, such as reducing belief in negative thoughts and inspiring solutions, compared to human-written stories from past DMH interventions?
To address these research questions, we conducted a randomized experiment with two groups of young adult participants: crowdworkers ($N=174$) and students ($N=172$). All participants engaged in a survey activity where they were randomly assigned to interact with either a human-written story from past interventions (Non-LLM story) or an LLM-enhanced story. Our findings revealed that LLM-enhanced stories were perceived as comparably authentic to Non-LLM stories, yet they performed better on average in communicating key takeaways, promoting reflection, reducing belief in negative thoughts, and inspiring actionable solutions. Our results suggest the promise of LLMs to create custom narratives supporting personal wellbeing, and inspire design considerations for LLM-enhanced narrative interventions, such as the need for a careful balance between relatability and avoiding implausibility and the importance of refining the tone and wording in AI-generated content.
@@ 2. Related Work
For the review of related work, we begin by discussing how story-based interventions can promote psychological wellbeing. Following this, we talk about the need for dynamic content generation in DMH interventions, focusing on the challenges and potential of LLMs in this context.
@@ 2.1. Promoting Psychological Wellbeing Through Stories
There is growing evidence supporting the benefits of narrative approaches in promoting psychological wellbeing and behavioral change (Bhattacharjee et al., 2022; Saksono et al., 2021; Grimes et al., 2008). Research has shown that storytelling interventions can reduce symptoms of depression across various populations (Vromans and Schweitzer, 2011; Rodríguez Vega et al., 2014); for instance, Rodríguez Vega et al. (2014) have utilized mindfulness-based narrative therapy as a therapeutic intervention for treating depression in cancer patients. Behavioral change interventions have leveraged stories to subtly guide audiences toward specific viewpoints or actions by immersing them in relatable narratives (Bilandzic and Busselle, 2013). Some interventions embed educational content within an entertainment framework containing narratives, engaging audiences while delivering important messages (Singhal and Rogers, 2012). Stories have been used successfully to encourage physical activity (Saksono et al., 2021), promote healthy eating habits (Grimes et al., 2008), and counter stigma against mental health conditions (Hospital, 2022; Evans-Lacko et al., 2014; Shoes, 2022) and sexual harassment victims (Dimond et al., 2013). This broad applicability underscores the power of stories to inspire behavioral change and foster psychological wellbeing in diverse contexts.
2.2. Dynamic Content Generation in DMH Interventions
Because mental health challenges are deeply personal and varied, static or pre-generated DMH content often fails to resonate with users (Bhattacharjee et al., 2024a). Dynamic content generation addresses this shortcoming by producing personalized and contextually relevant outputs (Datta et al., 2004). In the context of DMH interventions, this capability is critical for enhancing engagement, effectiveness, and user satisfaction (Bhattacharjee et al., 2023b; Baumel et al., 2019). However, delivering these benefits remains challenging.
2.2.1. Challenges of Dynamic Content Generation:
@@
Some prior studies (Ma et al., 2023; Morris and Picard, 2014; Smith et al., 2021; Saksono et al., 2023; Morris et al., 2015) explored supporting user wellbeing through dynamic content generated via crowdsourcing. For instance, Panoply (Morris and Picard, 2014; Morris et al., 2015), a crowd-powered system, allowed users to submit descriptions of their stressors. These descriptions were then sent to crowdworkers who crafted supportive messages in response. Similarly, Flip*Doubt was a crowd-powered web application that provided users with cognitive reappraisals of negative thoughts (Smith et al., 2021). Additionally, many online peer support groups have been established with a similar motivation (Baumel et al., 2018; Wang et al., 2023). These platforms aim to foster wellbeing by connecting individuals facing similar challenges, enabling them to share stories, learn from each other, and engage in discussions that promote reflection. However, these platforms heavily depend on the availability and responsiveness of participants, which is not always guaranteed. Moreover, the varying levels of comfort among crowdworkers and peers in sharing their experiences could limit the effectiveness of the support provided. Furthermore, crowdworkers and peers may also require training to ensure they deliver appropriate and empathetic support (Chen et al., 2021).
@@
The advent of machine learning and deep learning brought some advancements in dynamic content generation for DMH interventions. These models began to harness personalized data, such as user history, mood patterns, and behavioral cues, to generate responses that were more attuned to each user (Torous et al., 2018; Adler et al., 2021; Fitzpatrick et al., 2017). Numerous conversational chatbots were developed during this era, using AI algorithms to choose from pre-set branches (Fitzpatrick et al., 2017; Meyerhoff et al., 2024; Aguilera et al., 2020), but their capacity to process open-ended input from users and adapt accordingly was still very limited. Natural language processing (NLP) techniques, like information retrieval or word embedding (Carpenter et al., 2016), played important role in improving how systems interpreted free-form user input, enabling more relevant and context-sensitive outputs (Le Glaz et al., 2021). For instance, Morris et al. (2018) have employed these techniques to repurpose responses from existing online peer support databases, presenting the content that most align with their specific concerns. However, despite these advancements, the ability to dynamically generate new content still remained limited. These systems excelled at pattern recognition but fell short in the creativity needed to produce truly customized content without large-scale training datasets. As a result, most systems continued to rely on selecting responses from a predefined content bank rather than creating entirely new content (Abd-Alrazaq et al., 2019).
2.2.2. Potential of LLMs in Dynamic Content Generation for DMH Interventions
@@
Recent advancements in the development of LLMs like GPT and Gemini represent a leap forward for dynamic content generation. These models have the potential to generate coherent and rich text, offering a significant improvement over earlier systems that relied on selecting from predefined content (Bhattacharjee et al., 2024b). These capabilities can extend to various domains, particularly in DMH interventions, where understanding nuanced user needs is crucial. LLMs have the potential to tailor interventions to an individual’s unique experience that might involve a challenge they face or a recent experience. This is particularly facilitated by the LLMs’ ability to process open-ended input, allowing users to express and respond to their situations in rich detail and in any form they choose.
@@
For example, Sharma et al. (2023b) developed an LLM-based cognitive reframing tool that dynamically adapts its responses based on the specific challenges faced by users. Similarly, CLOVA CareCall, a conversational AI, collects data on individuals’ general health and engages as a conversational partner to alleviate loneliness by generating dynamic, human-like questions and responses (Jo et al., 2023, 2024). Additionally, Hedderich et al. (2024) introduced an LLM-based chatbot designed to assist teachers in providing support to adolescents dealing with cyberbullying, with responses that recognize and adapt to the context of each situation. Studies such as the one by Kim et al. (2024) have utilized LLMs to offer contextual writing support in journal writing. In their work, Kim et al.’s system provided real-time assistance as users expanded their journal entries. Collectively, these applications underscore the promise of LLMs in generating dynamic support that not only acknowledges but also responds appropriately to the nuanced needs of users.
@@
However, adapting LLMs for dynamic content generation in DMH tools is not a straightforward task. Research has shown that LLMs can produce inaccurate or clinically invalid suggestions within the wellbeing context (Ma et al., 2024; Demszky et al., 2023). This highlights the need for design considerations to ensure that content is not only accurate but also delivered effectively to support wellbeing. Additionally, it is crucial to determine which intervention elements can be personalized and which must remain standardized. Bhattacharjee et al. (2024b) emphasize that LLMs should operate within specific evidence-based psychological strategies when providing support in managing emotions. In this domain, LLMs require careful direction regarding the situational context of the user and should adhere strictly to predetermined structures and content guidelines.

### v4

# Perfectly to a Tee: Understanding User Perceptions of Personalized LLM-Enhanced Narrative Interventions

@@ Abstract.
Stories about overcoming personal struggles can effectively illustrate the application of psychological theories in real life, yet they may fail to resonate with individuals’ experiences. In this work, we employ large language models (LLMs) to create tailored narratives that acknowledge and address unique challenging thoughts and situations faced by individuals. Our study, involving 346 young adults across two settings, demonstrates that personalized LLM-enhanced stories were perceived to be better than human-written ones in conveying key takeaways, promoting reflection, and reducing belief in negative thoughts. These stories were not only seen as more relatable but also similarly authentic to human-written ones, highlighting the potential of LLMs in helping young adults manage their struggles. The findings of this work provide crucial design considerations for future narrative-based digital mental health interventions, such as the need to maintain relatability without veering into implausibility and refining the wording and tone of AI-enhanced content.
@@
journalyear: 2025copyright: acmlicensedconference: Designing Interactive Systems Conference; July 5–9, 2025; Funchal, Portugalbooktitle: Designing Interactive Systems Conference (DIS ’25), July 5–9, 2025, Funchal, Portugaldoi: 10.1145/3715336.3735810isbn: 979-8-4007-1485-6/2025/07ccs: Human-centered computing Empirical studies in HCI
Figure 1. Process of Generating Personalized LLM-Enhanced Narratives: User inputs their circumstances through structured and open-ended responses, which are combined with narrative guidelines and examples from past interventions, allowing the LLM to generate unique, personalized stories for each user.
[…]
Process of Generating Personalized LLM-Enhanced Narratives: User inputs their circumstances through structured and open-ended responses, which are combined with narrative guidelines and examples from past interventions, allowing the LLM to generate unique, personalized stories for each user.
@@ 1. Introduction
Stories about individuals overcoming personal struggles can serve as a powerful medium for promoting wellbeing (Bhattacharjee et al., 2022; Anderson and Wallace, 2015; Llewellyn-Beardsley et al., 2019; Kornfield et al., 2022b). When individuals engage with narratives that reflect their own experiences, they can draw parallels between their lives and those of the characters, creating opportunities for self-reflection (Ferguson and Musheno, 2000). For instance, a narrative about someone overcoming disappointment from job search rejection could provide valuable insights to others facing similar setbacks. This reflection process helps readers identify patterns in their thoughts or behaviors, leading to practical insights that empower them to make positive changes (Bhattacharjee et al., 2023a, 2022). Furthermore, engaging with such stories can help validate and normalize mental health challenges (Kornfield et al., 2022a). Given these substantial benefits, stories are increasingly being integrated into digital mental health (DMH) interventions (Meyerhoff et al., 2024; Bhattacharjee et al., 2022; Baumel et al., 2018).
@@
However, many DMH interventions face significant challenges in delivering engaging and applicable content (Hornstein et al., 2023; Bhattacharjee et al., 2023b). A key issue is the lack of tailoring to users’ unique struggles (Matthews and Rhodes-Maquaire, 2024), which can make the content feel generic and disconnected from their immediate needs. Such perceptions not only increase the likelihood that the content will be overlooked but also risk dismissing users’ genuine struggles (Bhattacharjee et al., 2023b). Consequently, seemingly irrelevant narratives may undermine the potential effectiveness of a DMH intervention (Bhattacharjee et al., 2022). Although some interventions have attempted to address this issue by using pre-defined rules for basic content adjustments, these strategies are inherently limited in the complexity of inputs they can handle and the range of personalization that can be applied during live interactions.
@@
Large language models (LLMs) offer promise in addressing these challenges by leveraging their ability to process vast amounts of data and adapt to diverse user inputs. They are rapidly gaining recognition for their capabilities as personalized intervention tools and supportive scaffolds in a variety of settings (Bhattacharjee et al., 2024c; Kim et al., 2023; Abd-Alrazaq et al., 2023). Given their flexibility, LLMs can interpret users’ descriptions of their situations in order to create stories specifically tailored to their circumstances. We posit that such capability of dynamic personalization – adapting content in real-time based on user inputs (Mahmood et al., 2014) – will significantly enhance the relevance and impact of digital narrative interventions. Revisiting the previous example of job search rejection, a user might gain more benefits from a narrative tailored to reflect the competitive nature of their field or environment. By connecting the story to specific circumstances and concerns, a narrative might provide deeper insights and more appropriate strategies, effectively addressing the user’s comprehensive experience.
@@
Although LLMs show potential for generating personalized stories, they are not inherently trained to adhere to established guidelines or structures (Bhattacharjee et al., 2024c). Given the sensitivity of mental health challenges, implementing an LLM-driven intervention requires careful measures to prevent generating harmful or insensitive content (Demszky et al., 2023). To address these risks, LLMs can be preloaded with narratives and mental health content from prior DMH interventions, constraining them to generate content within those boundaries. This approach would allow personalization to adapt the story’s context to match an individual’s circumstances while ensuring the core takeaways and guidance remain consistent with validated methods.
In this work, we developed an intervention designed to dynamically personalize stories to support individuals experiencing challenging thoughts. The intervention was designed to ensure that each participant received a story uniquely tailored to their specific circumstances. As shown in Figure 1, the intervention achieves dynamic personalization by incorporating user input into the prompt, where participants describe their challenging thoughts and provide open-ended descriptions of the difficulties they are facing. The LLM is also prompted to follow narrative intervention guidelines to ensure the generated story aligns with the intended format and qualities. Our goal was to understand how well this personalization approach can adhere to the intended qualities of the narratives while also assessing their impact as an intervention on specific outcomes, such as reducing belief in challenging thoughts.
@@
Our research focused on young adults aged 18–25, a demographic particularly susceptible to mental health issues due to the inherent transitional nature of this life stage, encompassing significant changes in education, career, and personal relationships (Kornfield et al., 2022a; Park, 2018). Statistically, this age group exhibits a much higher prevalence of mental health concerns compared to other adult groups (National Institute of Mental Health (2020), NIMH). These factors make them a relevant group for our research, where we aim to design stories that can help manage common challenges they face.
@@
  • RQ1: How do young adults perceive the key narrative qualities of personalized LLM-enhanced stories, such as authenticity and the ability to communicate takeaways, compared to human-written stories from past DMH interventions?
  • RQ2: How effective are personalized LLM-enhanced stories as interventions for managing negative thoughts, such as reducing belief in negative thoughts and inspiring solutions, compared to human-written stories from past DMH interventions?
To address these research questions, we conducted a randomized experiment with two groups of young adult participants: crowdworkers ($N=174$) and students ($N=172$). The intervention was implemented as a single-session intervention (SSI) (Schleider and Weisz, 2017), where participants received and responded to a single story. Participants were randomly assigned to engage with either a human-written story from past interventions (Non-LLM story) (Meyerhoff et al., 2024) or a personalized LLM-enhanced story (LLM story). We evaluated the impact that LLM stories had on participants’ beliefs in their negative thoughts and associated emotions relative to Non-LLM stories. We also gathered feedback on participants’ perceptions of the narratives, focusing on attributes such as the authenticity of the stories and the clarity with which key takeaways were communicated. Our findings revealed that LLM stories were perceived as comparably authentic to Non-LLM stories, yet they performed better on average in communicating key takeaways, promoting reflection, reducing belief in negative thoughts, and inspiring actionable solutions. Our results suggest the promise of LLMs to create custom narratives supporting personal wellbeing, and inspire design considerations for personalized LLM-enhanced narrative interventions, such as the need for a careful balance between relatability and avoiding implausibility and the importance of refining the tone and wording in AI-generated content.
Our contributions include:
  • Design of an intervention for delivering personalized LLM-enhanced narratives that align with users’ challenging thoughts and circumstances.
  • A randomized experiment conducted in two distinct settings, comparing personalized LLM-enhanced narratives with narratives from previous DMH interventions.
  • Design implications to guide future DMH interventions in utilizing LLMs for personalizing narrative-based support.
@@ 2. Related Work
For the review of related work, we begin by discussing how story-based interventions can promote psychological wellbeing. Following this, we talk about the need for dynamic personalization in DMH interventions, focusing on the challenges and potential of LLMs in this context.
@@ 2.1. Promoting Psychological Wellbeing Through Stories
There is growing evidence supporting the benefits of narrative approaches in promoting psychological wellbeing and behavioral change (Bhattacharjee et al., 2022; Saksono et al., 2021; Grimes et al., 2008; Park et al., 2021; Collins et al., 2022; Davis et al., 2020). Research has shown that storytelling interventions can reduce symptoms of depression across various populations (Vromans and Schweitzer, 2011; Rodríguez Vega et al., 2014); for instance, Rodríguez Vega et al. (2014) have utilized mindfulness-based narrative therapy as a therapeutic intervention for treating depression in cancer patients. Behavioral change interventions have leveraged stories to subtly guide audiences toward specific viewpoints or actions by immersing them in relatable narratives (Bilandzic and Busselle, 2013). Some interventions embed educational content within an entertainment framework containing narratives, engaging audiences while delivering important messages (Singhal and Rogers, 2012). Stories have been used successfully to encourage physical activity (Saksono et al., 2021), promote healthy eating habits (Grimes et al., 2008), and counter stigma against mental health conditions (Hospital, 2022; Evans-Lacko et al., 2014; Shoes, 2022) and sexual harassment victims (Dimond et al., 2013). This broad applicability underscores the power of stories to inspire behavioral change and foster psychological wellbeing in diverse contexts.
Narrative interventions can be delivered in multiple formats. In our work, we utilize them as SSIs (Schleider et al., 2020; Schleider and Weisz, 2017) — structured, standalone sessions designed to provide focused support that has immediate effects on specific outcomes. While sustained engagement over a longitudinal intervention can have enhanced benefits, we chose this format for our exploratory study since our research questions center around the impact of LLM integration. The intervention was designed to prompt participants to reflect on their challenging thoughts by reading a story about someone experiencing a similar struggle and making progress, encouraging them to reevaluate their beliefs. Similar to prior research that has evaluated the short-term benefits of SSIs (Sharma et al., 2023b; Bhattacharjee et al., 2024a), we evaluate the utility of our intervention by examining the immediate impact it has on participants’ belief in their challenging thoughts and the associated emotional intensity.
2.2. Dynamic Personalization in DMH Interventions
Because mental health challenges are deeply personal and varied, static or pre-generated DMH content often fails to resonate with users (Bhattacharjee et al., 2024a). The ability to dynamically personalize content in real time based on users’ specific input and evolving needs has therefore become a highly sought-after feature, as users often prefer intervention content that reflects their specific challenges (Jahedi et al., 2024; Garrigós et al., 2003). In the context of DMH, real-time personalization can enhance user engagement, increase intervention effectiveness, and improve overall user satisfaction (Bhattacharjee et al., 2023b; Baumel et al., 2019). Attaining such benefits not only requires an intervention to be able to process and respond to user input, but also to understand how to change the original content without deviating too far from its original materials. As described below, implementing these affordances remains a significant challenge.
2.2.1. Challenges of Dynamic Personalization:
@@
Some prior studies (Ma et al., 2023; Morris and Picard, 2014; Smith et al., 2021; Saksono et al., 2023; Morris et al., 2015) explored supporting user wellbeing through dynamic content generated via crowdsourcing. For instance, Panoply (Morris and Picard, 2014; Morris et al., 2015), a crowd-powered system, allowed users to submit descriptions of their stressors. These descriptions were then sent to crowdworkers who crafted supportive messages in response. Similarly, Flip*Doubt was a crowd-powered web application that provided users with cognitive reappraisals of negative thoughts (Smith et al., 2021). Additionally, many online peer support groups have been established with a similar motivation (Gatos et al., 2021; Baumel et al., 2018; Wang et al., 2023; Collins et al., 2022). These platforms aim to foster wellbeing by connecting individuals facing similar challenges, enabling them to share stories, learn from each other, and engage in discussions that promote reflection. However, these platforms heavily depend on the availability and responsiveness of participants, which is not always guaranteed. Moreover, the varying levels of comfort among crowdworkers and peers in sharing their experiences could limit the effectiveness of the support provided. Furthermore, crowdworkers and peers may also require training to ensure they deliver appropriate and empathetic support (Chen et al., 2021).
@@
The advent of machine learning and deep learning brought some advancements in dynamic personalization for DMH interventions. These models began to harness personalized data, such as user history, mood patterns, and behavioral cues, to generate responses that were more attuned to each user (Torous et al., 2018; Adler et al., 2021; Fitzpatrick et al., 2017). Numerous conversational chatbots were developed during this era, using AI algorithms to choose from pre-set branches (Fitzpatrick et al., 2017; Meyerhoff et al., 2024; Aguilera et al., 2020), but their capacity to process open-ended input from users and adapt accordingly was still very limited. Natural language processing (NLP) techniques, like information retrieval or word embedding (Carpenter et al., 2016), played important roles in improving how systems interpreted free-form user input, enabling more relevant and context-sensitive outputs (Le Glaz et al., 2021). For instance, Morris et al. (2018) have employed these techniques to repurpose responses from existing online peer support databases, presenting the content that most align with their specific concerns. However, despite these advancements, the ability to dynamically generate new content still remained limited. These systems excelled at pattern recognition but fell short in the creativity needed to produce truly customized content without large-scale training datasets. As a result, most systems continued to rely on selecting responses from a predefined content bank rather than creating entirely new content (Abd-Alrazaq et al., 2019).
2.2.2. Potential of LLMs in Dynamic Personalization for DMH Interventions
@@
Recent advancements in the development of LLMs like GPT and Gemini represent a leap forward for dynamic personalization. These models have the potential to dynamically generate coherent and rich text, offering a significant improvement over earlier systems that relied on selecting from predefined content (Bhattacharjee et al., 2024c). These capabilities can extend to various domains, particularly in DMH interventions, where understanding nuanced user needs is crucial. LLMs have the potential to tailor interventions to an individual’s unique experience that might involve a challenge they face or a recent experience. This is particularly facilitated by the LLMs’ ability to process open-ended input, allowing users to express and respond to their situations in rich detail and in any form they choose.
@@
For example, Sharma et al. (2023b) developed an LLM-based cognitive reframing tool that dynamically adapts its responses based on the specific challenges faced by users. Similarly, CLOVA CareCall, a conversational AI, collects data on individuals’ general health and engages as a conversational partner to alleviate loneliness by generating dynamic, human-like questions and responses (Jo et al., 2023, 2024). Additionally, Hedderich et al. (2024) introduced an LLM-based chatbot designed to assist teachers in providing support to adolescents dealing with cyberbullying, with responses that recognize and adapt to the context of each situation. Studies such as the one by Kim et al. (2024) have utilized LLMs to offer contextual writing support in journal writing. In their work, Kim et al.’s system provided real-time assistance as users expanded their journal entries. Collectively, these applications underscore the promise of LLMs in enabling dynamic personalization that not only acknowledges but also responds appropriately to the nuanced needs of users.
@@
However, adapting LLMs for dynamic personalization in DMH tools is not a straightforward task. Research has shown that LLMs can produce inaccurate or clinically invalid suggestions within the wellbeing context (Ma et al., 2024; Demszky et al., 2023). Therefore, interventions must ensure that content is not only accurate but also delivered effectively to support wellbeing. Additionally, it is crucial to determine which intervention elements can be personalized and which must remain standardized. Bhattacharjee et al. (2024c) emphasize that LLMs should operate within specific evidence-based psychological strategies when providing support in managing emotions. In this domain, LLMs require careful direction regarding the situational context of the user and should adhere strictly to predetermined structures and content guidelines.
In the context of our study, we designed our intervention to maintain desirable qualities such as authenticity and the ability to communicate key takeaways. We achieve this by not only providing guidance in the prompt, but also by feeding stories from past DMH interventions into the LLMs. We analyzed the quality of personalization by examining how the personalized stories preserved these desirable narrative qualities.